# PATCH: a deep learning method to assess heterogeneity of artistic practice in historical paintings

### v2

@@ Discussion
To our knowledge, ours is the first study to demonstrate this learning method and apply it both to paintings from a controlled experiment and to historical paintings by a well-known artist with a workshop. Among AI/ML-based methods designed to aid in the analysis of historical paintings, PATCH is notable in that it requires no external, ground truth data. Networks trained on known examples can accomplish tasks associated with workshop practice and authorship with exceptional accuracy. For example, Ugail and colleagues recently used transfer learning to authenticate paintings by the Renaissance master Raphael (Raffello Sanzio)34. By combining edge detection with typical feature extraction in a residual neural network, they were able to recognize works by Raphael with 98% accuracy. However, known examples of artists’ work may not be available, as exemplified by the workshop context. Our method shows exceptional promise as a tool for ML analysis of complex historical works where known samples of work by the artists involved do not exist. Indeed, it may have extensive application in other image analysis tasks where there is little to no ground truth information available.
Further, the second phase of PATCH creates what we consider a measure of the heterogeneity of artistic practice (HAP) within a given dataset. Communities found within a dataset represent different artists, materials or both. The degree to which those communities are interconnected provides information about the possibility of shared authorship, materials, or other practices. A fully connected network with no distinct communities has . As the number of communities increases and the interconnection between communities decreases, modularity increases toward 1 (a completely disconnected network has undefined modularity, but would effectively represent the maximum heterogeneity). The networks from our three studies provide empirical examples of lower and higher modularity networks. Christ on the Cross and The Baptism have values that cluster around 0.3, considered the threshold for evidence of structure, indicating low heterogeneity. The network for the experimental student paintings has much greater heterogeneity at (see Figure S3). The maximum modularity depends on the number of edges in the network35, but also on the number of communities in the network. A network with two fully connected and completely disjoint communities has , for example. As such, network modularity can function as a measure of absolute observed heterogeneity (Figure 6).
@@
At first glance, the communities identified in our analysis of El Greco’s Baptism appear to roughly correspond to the proposed attributions by previous authors discussed above. The community in yellow in Figure 3 overlaps much of the top of the painting, previously attributed to El Greco, while the community in blue overlaps John the Baptist and the figure in red, assigned to Jorge Manuel. We also find a combination of communities in the face and torso of Christ, which was attributed to both artists. A single, small and particularly unique community (in orange in Figure 3) overlaps the landscape at the bottom of the painting and Christ’s legs.
@@
Overall, however, the PATCH findings suggest a different story underlying the existing art historical interpretation. The partition of The Baptism has , which is only slightly above the threshold for evidence of structure. Nearly a third of all edges in the network are interconnections between three of the four communities we have identified. The yellow and green groups have 49 intercommunity edges, blue and green 30, and blue and yellow 22. If the blue, green and yellow communities indeed represented three distinct artists working with the same or similar materials, we would expect the communities to be more insular and the modularity of the network to be substantially higher, similar to what we observed in our controlled test. Yet the modularity of The Baptism is much closer to that of Christ on the Cross (), which appears to feature one artist. It is worth noting that, as shown in the analysis of student paintings, the CNN is particularly proficient at recognizing different-artist pairs and is more likely to incorrectly label a same-artist pair as different than the converse. Thus, the intercommunity edges in The Baptism network that survived our initial pruning are not liable to be spurious.
This suggests that some unifying factor connects these three communities. One possibility is a single artist working with different brushes. In El Greco’s time, brushes were made with either bristles (which are coarse) or hair (which is fine and soft). Each served distinct technical purposes and often came in material-specific shapes36. Variation could arise from using the hair/bristles of different species in the same type of brush. Key sources were hog bristles (especially white hog) and “minever” (gray squirrel tail fur)37, though dog and mongoose often replaced minever due to its rarity36,38. Hog’s bristles likewise may have been replaced with ox39. Such changes could affect paint application. For example, hog’s bristles are naturally flagged (the ends of the hairs are split), allowing them to pick up more paint40. Bristle substitution could decrease brush capacity. Further experimentation to understand the effect of materials on PATCH analysis is needed.
@@
The evidence could also indicate the work of one artist with changing style or technique over time. Interestingly, Biannuci and colleagues41 present evidence that El Greco suffered a series of ischemic events, one in the 1590s and a second in 1608 (the year the contract for the altarpieces for the Hospital Tavera was initiated), “resulting in progressive disabilities with fluctuating course characterized by temporary improvements and worsening before his death.” The regions that were previously identified as the work of different individual painters could conceivably represent variation in the master’s individual style (however this may have arisen) over the course of his final years. Investigation of changes in style and technique and their effect on PATCH analysis will be an important direction for future research, as well.
PATCH has the capability to make a substantial contribution to research as a complement to existing art historical methods. The accuracy of the method has been demonstrated in the validation on student paintings. Our analyses of Christ on the Cross and The Baptism show the true potential of PATCH by contributing important new information to art historical scholarship on the creation of objects in El Greco’s workshop in Toledo.
@@
In the analysis of The Baptism, the fact that the orange group, which overlaps a region known to have been altered after the delivery of the painting, is so disconnected from the remainder of the communities speaks to the method’s capability of recognizing a unique contribution or substantial departure from the technique used in the rest of the painting. That the other three communities roughly correspond to art historical attributions suggests PATCH possesses a kind of semantic fidelity42—our method recognizes areas that appear different to the trained eye of an art historian. However, the connections between those areas revealed by PATCH call this attribution into question. Traditional attribution studies rely on the inspection of painters’ unconscious tendencies and individual styles. For example, in attributing the painting of John the Baptist to El Greco’s son, Wethey29 cites the “hardness” of the details. Computer vision allows us to view the surface of a painting at a different scale. Previous analyses indicate that features as small as the diameter of a single paint brush bristle (length scales of 0.2–0.4 mm) may be integral to the network’s identification of an artist17. By analyzing spatial correlations at this microscopic scale, the network may be revealing aspects of the physics of paint application and of the physiology of hand movements. PATCH adds micro-scale features to the macro (brushstrokes) and meta-scale (historical and material) data, allowing for a full-scale analysis and the discovery of heretofore unseen evidence of the processes by which paintings were created.
@@
Future research with PATCH will focus on teasing apart the influences of materials and artists on the heterogeneity of artistic practice. Expansion of PATCH applications to include inter-painting comparisons and other art historical applications will be important as well. However, uses of PATCH are not limited to those contexts: it provides a general-purpose approach to clustering data that can be encoded in the form of image-like arrays, as occurs in a wide variety of applications across the social and natural sciences. Underlying the effectiveness of pairwise assignment training is the power of convolutional neural networks to find extremely subtle differences among such data points, and the counter-intuitively useful information gained when the network consistently fails to succeed in this task. As such, PATCH can be deployed where supervised learning is impossible (because the ground truth is unknown) and when alternative unsupervised approaches are ineffective.
@@ Acknowledgments
The authors would like to thank Fundación Casa Ducal de Medinaceli for generously providing data for The Baptism, and Sr. Juan Manuel Albendea Solis for his insightful criticism that helped us reevaluate our conclusions and improve the manuscript. We would also like to thank the Cleveland Museum of Art for use of the data from Christ on the Cross, and the Lapis Senior Conservator of Paintings and Head of Paintings Conservation, Dean Yoder, for his help and feedback. The authors gratefully acknowledge support from the Expanding Horizons Initiative of Case Western Research University (CWRU) College of Arts and Sciences and the Jack, Joseph and Morton Mandel Foundation. Computational support was provided by the CWRU High Performance Computing cluster and a workstation kindly donated through the Dell Seed Program.
@@
Appendix B: Hyperparameter Optimization
@@
Appendix C: Comparison to alternative methods
@@
Pairwise comparison of large numbers of regions via ML is computationally expensive, so a simpler method would be preferable if it were equally accurate. One simple way of distinguishing between regions of a painting using topography would be to compare the “roughness” of the surfaces. We applied a roughness-based method to the student paintings to determine its accuracy relative to the PATCH method. We used the standard deviation of height values as a measure of roughness, and calculated the standard deviation of height values for each 1x1 cm2 patch of each student painting . We then compared paintings pairwise using the pairwise Wilcoxon rank sum test implemented in R using the stats package. A significant result on the pairwise Wilcoxon test means the distributions (specifically the medians) are significantly different, which in this context would indicate a different-artist pair. Thus, same-artist pairs would be tests where the null hypothesis could not be rejected. Using the Bonferroni correction for multiple tests with =0.01, any pairwise comparison with a p-value greater than 0.00003 was considered a same-artist pair. This method had an overall accuracy of 81.0%, identifying 19 of 23 same-artist pairs (82.6%) and 220 of 277 different artist pairs (79.4%), a significant reduction in accuracy relative to PATCH. This translates to F1 scores of 0.384 for same-artist pairs and 0.878 for different-artist pairs, an average of 0.631 (Table S2). We conclude that PATCH substantially outperforms roughness for identification of painters. This also indicates that the network is not simply sorting patches of images based on their roughness.
| Precision | Recall | F1 | |
| Same-artist pairs | 0.250 | 0.826 | 0.384 |
| Different-artist pairs | 0.982 | 0.794 | 0.878 |
| Average | 0.616 | 0.810 | 0.631 |
Unsupervised method
@@
Feature extraction: each patch was converted to a feature vector using a 2D Fourier transform, which created a Fourier domain 200x200 patch with the origin in the center. We took the absolute values of the upper half of the domain (the bottom was identical by symmetry), leaving out the 13 central columns. These columns correspond to the largest wavelengths (comparable to the patch size), and deemphasizing them puts the focus on brushstroke details at smaller wavelengths. We flattened the absolute values into a vector of size 17578. Other parameter choices gave comparable or worse final performance. We also tried an alternative feature extractor based on the VGG16 convolutional neural network (CNN) that had been trained to classify height data in our earlier student painting controlled experiment2, with the bottom layers of the network cut out. This performed worse than the Fourier feature extractor, so we stayed with the latter.
@@
Dimension reduction: to allow eventual clustering of the feature vectors, we used the UMAP dimension reduction algorithm3, with default parameters, to map the full feature vectors onto 4-dimensional vectors.
@@
Clustering: Finally, we used the flat clustering module of HDBSCAN4 (minimum cluster size = 25, cluster selection method = ‘leaf’) to force the algorithm to group the 4-dimensional vectors into two clusters.
@@
If there was a statistically significant difference in the numbers of patches assigned to each cluster among the two paintings (chi-squared test with p 0.05), then the pair of paintings were marked as different artist. For example, if the labeled patches in one painting were mostly assigned to cluster 1, and the labeled patches in the other painting to cluster 2, this would be taken as evidence that different artists were involved. Otherwise (i.e. if cluster assignments were fairly evenly distributed among the two paintings) the pair was marked as same artist. The algorithm occasionally failed to find two clusters at all, which we also treated as evidence that the pair was made by the same artist. The final results were quite similar to the surface roughness approach described above: the method correctly identified 18 out of 23 same-artist pairs (78.3%), and 225 out of 277 different-artist pairs (81.2%), yielding F1 scores of 0.387 and 0.888 for the same-artist and different-artist cases respectively. Changing the HDBSCAN parameters like minimum cluster size and the cluster selection method did not improve the results, and in all cases the approach performed significantly worse than the PATCH method.
| Precision | Recall | F1 | |
| Same-artist pairs | 0.257 | 0.783 | 0.387 |
| Different-artist pairs | 0.978 | 0.812 | 0.888 |
| Average | 0.618 | 0.797 | 0.637 |
Appendix D: Additional Figures
Supplementary References
- [1] Bergstra, J., Komer, B., Eliasmith, C., Yamins, D. & Cox, D. D. Hyperopt: a Python library for model selection and hyperparameter optimization. Comput. Sci. Discov. 8, 014008 (2015).
- [2] Ji, F. et al. Discerning the painter’s hand: machine learning on surface topography. Heritage Science 9, 152 (2021).
- [3] McInnes, L., Healy, J., Saul, N. & Großberger, L. UMAP: Uniform Manifold Approximation and Projection. Journal of Open Source Software 3, 861 (2018).
- [4] McInnes, L., Healy, J. & Astels, S. HDBSCAN: Hierarchical density based clustering. Journal of Open Source Software 2, 205 (2017).

### v3

@@ Discussion
To our knowledge, ours is the first study to demonstrate the PATCH learning method and apply it both to paintings from a controlled experiment and to historical paintings by a well-known artist with a workshop. Among AI/ML-based methods designed to aid in the analysis of historical paintings, PATCH is notable in that it requires no external, ground truth data. Networks trained on known examples can accomplish tasks associated with workshop practice and authorship with exceptional accuracy. For example, Ugail and colleagues recently used transfer learning to authenticate paintings by the Renaissance master Raphael (Raffello Sanzio)37. By combining edge detection with typical feature extraction in a residual neural network, they were able to recognize works by Raphael with 98% accuracy. However, known examples of artists’ work may not be available, as exemplified by the workshop context. Our method shows exceptional promise as a tool for ML analysis of complex historical works where known samples of work by the artists involved do not exist. Indeed, it may have extensive application in other image analysis tasks where there is little to no ground truth information available.
Further, the second phase of PATCH creates what we consider a measure of the heterogeneity of artistic practice (HAP) within a given dataset. Communities found within a dataset represent different artists, materials, or both. The degree to which those communities are interconnected provides information about the possibility of shared authorship, materials, or other practices. A fully connected network with no distinct communities has . As the number of communities increases and the interconnection between communities decreases, modularity increases toward 1 (a completely disconnected network has undefined modularity, but would effectively represent the maximum heterogeneity). The networks from our three studies provide empirical examples of lower and higher modularity networks. Christ on the Cross and The Baptism have values that cluster around 0.3, considered the threshold for evidence of structure, indicating low heterogeneity. The network for the experimental student paintings has much greater heterogeneity at (see Figure S4). The maximum modularity depends on the number of edges in the network38, but also on the number of communities in the network. A network with two fully connected and completely disjoint communities has , for example. As such, network modularity can function as a measure of absolute observed heterogeneity (Figure 5).
@@
At first glance, the communities identified in our analysis of The Baptism appear to roughly correspond to the proposed attributions by previous authors discussed above. The community in yellow in Figure 3 overlaps much of the top of the painting, previously attributed to El Greco, while the community in blue overlaps John the Baptist and the figure in red, assigned to Jorge Manuel. We also find a combination of communities in the face and torso of Christ, which was attributed to both artists. A single, small and particularly unique community (in orange in Figure 3) overlaps the landscape at the bottom of the painting and Christ’s legs.
@@
Overall, however, the PATCH findings suggest a different story underlying the existing art historical interpretation. The partition of The Baptism has , which is only slightly above the threshold for evidence of structure. Nearly a third of all edges in the network are interconnections between three of the four communities we have identified. The yellow and green groups have 49 intercommunity edges, blue and green 30, and blue and yellow 22. If the blue, green, and yellow communities indeed represented three distinct artists working with the same or similar materials, we would expect the communities to be more insular and the modularity of the network to be substantially higher, similar to what we observed in our controlled student painting dataset. Yet the modularity of The Baptism is much closer to that of Christ on the Cross (), which appears to feature one artist. It is worth noting that, as shown in the analysis of student paintings, the CNN is particularly proficient at recognizing different-artist pairs and is more likely to incorrectly label a same-artist pair as different than the converse. Thus, the intercommunity edges in The Baptism network that survived our initial pruning are not liable to be spurious.
@@
This suggests that some unifying factor connects these three communities. One possibility is a single artist working with different brushes. Renaissance artists used brushes made from the bristles or hair of several different mammal species, variation in the characteristics (e.g., coarseness) of which could affect the deposition of paint 39; 40; 41 (see Supplementary Materials for more information). The evidence could also indicate the work of one artist with changing style or technique over time. Interestingly, Biannuci and colleagues42 present evidence that El Greco suffered a series of ischemic events, one in the 1590s and a second in 1608 (the year the contract for the altarpieces for the Hospital Tavera was initiated), “resulting in progressive disabilities with fluctuating course characterized by temporary improvements and worsening before his death.” The regions that were previously identified as the work of different individual painters could conceivably represent variation in the master’s individual style (however this may have arisen) over the course of his final years. Investigation of changes in style and technique and their effect on PATCH analysis will be an important direction for future research.
@@
PATCH has the capability to make a substantial contribution to research as a complement to existing art historical methods. The accuracy of the method has been demonstrated in the validation on student paintings. Our analyses of Christ on the Cross and The Baptism show the true potential of PATCH by contributing important new information to art historical scholarship on the creation of objects in El Greco’s workshop in Toledo. In the analysis of The Baptism, the fact that the orange group, which overlaps a region known to have been altered after the delivery of the painting, is so disconnected from the remainder of the communities (see Supplementary Materials for further analyses) speaks to the method’s capability of recognizing a unique contribution or substantial departure from the technique used in the rest of the painting. That the other three communities roughly correspond to art historical attributions suggests PATCH possesses a kind of semantic fidelity43—our method recognizes areas that a trained art historian would classify as different. However, the connections between those areas revealed by PATCH also call those attributions into question. Computer vision allows us to view the surface of a painting at a different scale. Previous analyses indicate that features as small as the diameter of a single paint brush bristle (length scales of 0.2–0.4 mm) may be integral to the network’s identification of an artist2. By analyzing spatial correlations at this microscopic scale, the network may be revealing aspects of the physics of paint application and of the physiology of hand movements. PATCH adds micro-scale features to the macro (brushstrokes) and meta-scale (historical and material) data, allowing for a full-scale analysis and the discovery of heretofore unseen evidence of the processes by which paintings were created.
@@
Future research with PATCH will focus on teasing apart the influences of materials and artists on the heterogeneity of artistic practice. Expansion of PATCH applications to include intra-painting comparisons and other art historical applications will be important as well. However, uses of PATCH are not limited to those contexts: it provides a general-purpose approach to assessing regional similarity in images or data that can be encoded in image-like arrays. As such, PATCH is a promising method for applications where supervised learning is impossible (because the ”ground truth” is unknown) and alternative unsupervised approaches are ineffective, especially for a posteriori cluster number. Potential applications span the physical and social sciences, and include image segmentation tasks in applications such as medical imaging44, agricultural remote sensing6, urban development and design7, and microstructural analysis, as well as anomaly detection in a range of manufacturing contexts8; 9, among others.
Data availability
The dataset used to develop and validate the PATCH method is available on OSF at osf.io/a5xnh/. Height data from The Baptism and Christ on the Cross are the property of the institutions that own those paintings: the Fundación Casa Ducal de Medinaceli and the Cleveland Museum of Art, respectively.
Code availability
The code used to run the PATCH algorithm is available on GitHub: https://github.com/hincz-lab/PATCH.
@@ Acknowledgments
The authors would like to thank Fundación Casa Ducal de Medinaceli for providing data for The Baptism and Sr. Juan Manuel Albendea Solis for his feedback on our initial draft, which helped us reevaluate our conclusions and improve the manuscript. We would also like to thank the Cleveland Museum of Art for use of the data from Christ on the Cross, and the Lapis Senior Conservator of Paintings and Head of Paintings Conservation, Dean Yoder, for his help and feedback. The authors acknowledge support from the Expanding Horizons Initiative of Case Western Reserve University (CWRU) College of Arts and Sciences and the Jack, Joseph and Morton Mandel Foundation. Computational support was provided by the CWRU High Performance Computing cluster and a workstation donated through the Dell Seed Program.
Author contributions
AV conceived the PATCH method; LNS and MSM conceived the initial project; AV, MH, KS, MSM developed the method; AI oversaw creation of paintings; ALo, CB, and ITM performed scans; AV and MM wrote PATCH code; AV, ALi, MM and MH performed experiments and analyses; AV, MH, KS, ESB, CP, ALi, and LNS wrote the manuscript.
Ethics declarations
Competing interests: The authors declare no competing interests.
Supplementary information
Supplementary Information is available for this paper.
Correspondence
Correspondence and requests for materials should be addressed to AV.
@@
Appendix B: Hyperparameter Settings and Optimization
The first phase of PATCH analysis, pairwise assignment training, was performed using VGG16 architecture pretrained on the Imagenet database. Output from VGG-16 was pooled and flattened and passed through two dense layers with dropout rates of 0.25. The learning rate was set to 0.0001 and the batch size was set to 32 patches.
@@
Appendix C: PATCH in Comparison to Alternative Methods
The PATCH algorithm outperformed alternative methods (described below) for the student painting dataset, identifying same-artist pairs (n = 23) and different-artist pairs (n = 277) with excellent classification performance, summarized in Table S2. In the table, precision and recall are calculated by treating each category (same-artist pairs or different-artist pairs) as “positives” in turn. For example, for the same-artist pairs, the true positives TP number of same-artist pairs classified as same, false positives FP number of different-artist pairs classified as same, and false negatives FN number of same-artist pairs classified as different. The precision = , the recall = , and the F1 score is the harmonic mean of precision and recall. The definitions for the different-artist pair category are as above, but with “same” and “different” interchanged. The last three columns of the table show averages of the same-artist and different-artist columns for precision, recall, and F1. The PATCH method was superior to the tested alternatives in the majority of metrics, notably achieving the highest F1 scores for same-artist pairs and different-artist pairs. The closest competitor, supervised UMAP with a Fourier feature extractor, achieved only marginally higher precision for same-artist pairs and recall for different-artist pairs, beating PATCH by 2.9% and 0.3% respectively. However it performed substantially worse on same-artist recall (0.652 versus 0.87), thus scoring lower on F1. That method’s marginally increased precision is therefore outweighed by its inability to identify same-artist pairs (15/23, 6̃5%). In the following we describe the alternative methods we tested.
| Same-artist pairs () | Different-artist pairs () | Average | |||||||
| Method | Precision | Recall | F1 | Precision | Recall | F1 | Precision | Recall | F1 |
| PATCH | 0.909 | 0.87 | 0.889 | 0.989 | 0.993 | 0.991 | 0.949 | 0.931 | 0.940 |
| Surface roughness | 0.250 | 0.826 | 0.384 | 0.982 | 0.794 | 0.878 | 0.616 | 0.810 | 0.631 |
| Fourier / UMAP | 0.257 | 0.783 | 0.387 | 0.978 | 0.812 | 0.888 | 0.618 | 0.797 | 0.637 |
| Fourier / sUMAP | 0.8 | 0.522 | 0.632 | 0.961 | 0.989 | 0.975 | 0.881 | 0.755 | 0.803 |
| Fourier / UMAP | 0.254 | 0.652 | 0.366 | 0.967 | 0.841 | 0.9 | 0.611 | 0.747 | 0.633 |
| Fourier / sUMAP | 0.938 | 0.652 | 0.769 | 0.972 | 0.996 | 0.984 | 0.955 | 0.824 | 0.877 |
| Fourier / UMAP | 0.274 | 0.739 | 0.4 | 0.975 | 0.838 | 0.901 | 0.624 | 0.788 | 0.65 |
| Fourier / sUMAP | 0.688 | 0.478 | 0.564 | 0.958 | 0.982 | 0.97 | 0.823 | 0.73 | 0.767 |
| CNN / UMAP | 0.197 | 0.565 | 0.292 | 0.957 | 0.809 | 0.877 | 0.577 | 0.687 | 0.584 |
| CNN / sUMAP | 0.37 | 0.739 | 0.493 | 0.976 | 0.895 | 0.934 | 0.673 | 0.817 | 0.713 |
| CNN / UMAP | 0.23 | 0.609 | 0.333 | 0.962 | 0.83 | 0.891 | 0.596 | 0.72 | 0.612 |
| CNN / sUMAP | 0.34 | 0.739 | 0.466 | 0.976 | 0.881 | 0.926 | 0.658 | 0.81 | 0.696 |
| CNN / UMAP | 0.239 | 0.696 | 0.356 | 0.97 | 0.816 | 0.886 | 0.604 | 0.756 | 0.621 |
| CNN / sUMAP | 0.348 | 0.696 | 0.464 | 0.972 | 0.892 | 0.93 | 0.66 | 0.794 | 0.697 |
@@
Pairwise comparison of large numbers of regions via PATCH is computationally expensive, so a simpler method would be preferable if it were equally accurate. One simple way of distinguishing between regions of a painting using topography would be to compare the “roughness” of the surfaces. We applied a roughness-based method to the student paintings to determine its accuracy relative to the PATCH method. We used the standard deviation of height values as a measure of roughness, and calculated the standard deviation of height values for each 1x1 cm2 patch of each student painting. We then compared paintings pairwise using the pairwise Wilcoxon rank sum test implemented in R using the stats package. A significant result on the pairwise Wilcoxon test means the distributions (specifically the medians) are significantly different, which in this context would indicate a different-artist pair. Thus, same-artist pairs would be tests where the null hypothesis could not be rejected. Using the Bonferroni correction for multiple tests with =0.01, any pairwise comparison with a p-value greater than 0.00003 was considered a same-artist pair. This method resulted in F1 scores of 0.384 for same-artist pairs and 0.878 for different-artist pairs, substantially worse than PATCH. This also indicates that PATCH is not simply sorting patches of images based on their roughness, but using more complex characteristics of the height data for classification.
Unsupervised UMAP methods
@@
Feature extraction: each patch was converted to a feature vector using one of two feature extractors: i) a 2D Fourier transform, which created a Fourier domain 200x200 patch with the origin in the center. We took the absolute values of the upper half of the domain (the bottom was identical by symmetry), leaving out the 13 central columns. These columns correspond to the largest wavelengths (comparable to the patch size), and deemphasizing them puts the focus on brushstroke details at smaller wavelengths. We flattened the absolute values into a feature vector of size 17578. Other parameter choices gave comparable or worse final performance; ii) a feature extractor based on the VGG16 convolutional neural network (CNN) that had been trained to classify height data in our earlier student painting controlled experiment 2. The bottom layers of the network were cut out, and the output was pooled and flattened to yield a feature vector of size 512. These two types of feature extractors are labeled as “Fourier” or “CNN” respectively in Table S2.
@@
Dimension reduction: to allow eventual clustering of the feature vectors, we used the UMAP dimension reduction algorithm 46 with default parameters, to map the full feature vectors onto -dimensional vectors. Three different choices of were tested: , , and .
@@
Clustering: Finally, we used the flat clustering module of HDBSCAN 47 (minimum cluster size = 25, cluster selection method = ‘leaf’) to force the algorithm to group the -dimensional vectors into two clusters.
@@
If there was a statistically significant difference in the numbers of patches assigned to each cluster among the two paintings (chi-squared test with p 0.05), then the pair of paintings were marked as different artist. For example, if the labeled patches in one painting were mostly assigned to cluster 1, and the labeled patches in the other painting to cluster 2, this would be taken as evidence that different artists were involved. Otherwise (i.e. if cluster assignments were fairly evenly distributed among the two paintings) the pair was marked as same artist. The algorithm occasionally failed to find two clusters at all, which we also treated as evidence that the pair was made by the same artist. The resulting classification metrics, regardless of feature extractor or dimension , were quite similar to those of the surface roughness approach described above. Changing HDBSCAN parameters like minimum cluster size and the cluster selection method did not substantially alter the performance, and in all cases the unsupervised approach fared worse than the PATCH method.
Supervised UMAP (sUMAP) methods
The final set of methods we tested were supervised variants of the UMAP procedure above. These take advantage of the fact that UMAP can be nudged to provide more distinct separation of clusters by giving it class labels. The approach is the same as earlier, with three differences: 1) 70% of the patches in each of the two paintings are randomly selected to be training data. These patches are given distinct labels by painting (i.e. painting 1 versus painting 2), which are input to the UMAP algorithm along with the patches. 2) The remaining 30% of the patches are used to test whether the approach can successfully distinguish paintings by putting each painting in a distinct cluster. As before, if a statistically significant difference exists in the number of test patches assigned to each cluster among the two paintings (chi-squared test with ) then the pair is marked as different-artist. 3) Because of the stochastic nature of training / test split, the analysis was repeated three independent times, and where there was disagreement we used the majority outcome of the three runs to decide on same-artist or different-artist for each pair.
Both Fourier and CNN feature extractors were tested with this approach, as well as different output dimensions . The supervised component did improve the results, and as seen in Table S2 same-artist precision and different-artist recall were slightly better than PATCH (by 2.9% and 0.3% respectively). However the same-artist recall was relatively poor (0.652 for this method versus 0.87 for PATCH). Hence overall PATCH was still superior in classification, as evidenced by its higher F1 scores for both same-artist and different-artist pairs. Additionally, the computational complexity of this approach (unlike the simpler surface roughness and unsupervised UMAP methods) was comparable to implementing PATCH, so that there were no substantial performance benefits.
Appendix D: Analysis of PATCH-generated Communities
Further analysis of the community structure in the PATCH generated networks can provide additional insight into the community cohesiveness and distinction beyond the modularity index. These properties are quantified by the average community degree, broadly defined as the weighted and normalized sum of all edges between nodes within two community distinctions. This analysis can provide additional insights into the cohesiveness of artistic practice within a single community, and the relative similarity of practice between multiple communities.
The internal degree () is defined as the weighted sum of all edges () between nodes in the same community (), normalized by the theoretical maximum number of edges for available nodes:
The internal degree can be taken as a quantified metric of the intra-community cohesiveness, related to the variability of textures within an identified community. The external degree () between two communities is similarly defined as the weighted sum of all edges between nodes in distinct communities (, ), normalized by the theoretical maximum number of edges for and available nodes in the respective communities:
The external degree can be taken as a quantified metric for the similarity between two separate communities, related to the distinctiveness of textures between them. Note that due to the prior identification of distinct communities, the external degree is expected to be much lower than the internal. Finally, the total external degree for a community () is defined as the external degree between all nodes internal to () and external to () the community, again normalized by the theoretical maximum number of edges:
The total external degree is a quantified metric describing the relative isolation of a single community from the remainder of the network, related to the overall distinction of textures from the whole. In principle, both the pairwise and total degree metrics are useful in PATCH analysis. For instance, a community resulting from extensive conservation work could engender a low total external degree due to the difference in tools and artist purpose, while multiple communities from similar sources, such as multiple students of the same master, could show a similar high external degree describing similarity of practice. The internal degree could describe, for instance, the region-to-region consistency of a single artist, or the conformity of practice by multiple similar artist sources.
Figure S2 shows the computed internal and external network degrees for the PATCH analysis of the two El Greco paintings. Figure S2 A and C show a pairwise comparison between each of the identified communities in each painting, and Figure S2 B and D show the internal degree and total external degree for each community.
In The Baptism, the external degrees between the first three communities are significantly stronger than those involving the fourth community, with the average external degree over four times larger. This indicates a significant isolation of the fourth community (colored orange in Figure 3, and represented by orange bars in Figure S2 B). The strongest external degree is between communities 1 (yellow) and 3 (green) and the weakest is between 3 and 4, which has no connecting edges. Overall, the 3rd community has a slightly greater connection with the remainder of the network, indicating that the height map is the most similar to all other areas. The internal degree is similar for all communities (%), indicating that no community stands out as having a larger disparity in topographical signal. The external degree analysis is largely redundant in the case of Christ on the Cross, since only two communities are proposed. However, it should be noted that the external degree between the two communities (Christ) is significantly larger than for any other two communities in The Baptism except the connection between communities 1 and 3 (Baptism). This indicates that, on average, more similarities can be found between the two communities in Christ on the Cross than those proposed in The Baptism. This is reflected in the low modularity value. Additionally, the internal degree in each community is similar, which can indicate a similar intra-community stylistic distribution, and supports the two-community model.
Appendix E: Brush Construction in the Early Modern Period
In El Greco’s time, brushes were made with either bristles (which are coarse) or hair (which is fine and soft). Each served distinct technical purposes and often came in material-specific shapes 39. Variation could arise from using the hair/bristles of different species in the same type of brush. Key sources were hog bristles (especially white hog) and “minever” (gray squirrel tail fur) 41 , though dog and mongoose often replaced minever due to its rarity 39; 40. Hog’s bristles likewise may have been replaced with ox 48. Such changes could affect paint application. For example, hog’s bristles are naturally flagged (the ends of the hairs are split), allowing them to pick up more paint 49. Bristle substitution could decrease brush capacity. Further experimentation to understand the effect of materials on PATCH analysis is needed.
Appendix F: Additional Figures
Appendix G: Edge Pruning